# Can LLMs Hit Moving Targets? Tracking Evolving Signals in Corporate Disclosures

### v2

@@ Abstract
Moving targets (Cohen & Nguyen, 2024) – managers’ strategic shifting of key performance metrics when the original targets become difficult to achieve – have been shown to predict subsequent stock underperformance. However, our work reveals that the method employed in that study exhibits two key limitations that hinder the accuracy – noise in the extracted targets and loss of contextual information – both of which stem primarily from the use of a named entity recognition (NER). To address these two limitations, we propose an LLM-based target extraction method with a newly defined metric that better captures semantic context. This approach preserves semantic context beyond simple entity recognition and yields consistently higher predictive power than the original approach. Overall, our approach enhances the granularity and accuracy of financial text-based performance prediction.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have spurred efforts to leverage these tools for uncovering economically meaningful signals from unstructured financial texts such as earnings calls and corporate disclosures. In line with prior research, our work applies LLMs to enhance the target and metric extraction methodology proposed in Cohen & Nguyen (2024). Their study documents that managers strategically shift emphasis to new performance metrics when existing ones deteriorate – a “moving targets” behavior that predicts subsequent stock price declines.
@@ 2.1 Motivation
Prior work (Cohen & Nguyen, 2024) has relied heavily on NER to extract performance targets from earnings call transcripts. However, we found that this approach suffers from limitations in both accuracy and explainability.
@@ 2.2 LLM-Based Target Extraction
Semantic Similarity. In the original formulation of the moving targets (Cohen & Nguyen, 2024), the degree of movement is quantified as
@@
$MT_{c,t}=\frac{1}{n}\sum_{j=1}^{b}g(\tilde{S}_{j}).$
This semantic similarity–based metric extends the original discrete measure by preserving semantic equivalence even when targets are phrased differently. This metric quantifies the extent to which performance targets previously emphasized by a firm are dropped or semantically altered after four quarters, similar with Cohen & Nguyen (2024).
@@ 3.1 Data and Summary Statistics
We conduct our analysis using earnings call transcripts as the primary source of firm-level disclosures, following Cohen & Nguyen (2024). While we also experimented with 10-K and 10-Q MD&A sections, we found that earnings calls provided the most informative setting for detecting shifting in targets. Our study focuses on a restricted sample of S&P 100 firms due to computational cost. The sample covers between January 2010 and December 2024, producing 5,615 firm-quarter observations across 64 quarters.
@@ 3.2 Implementation Details
For validation, we adopt the empirical design of Cohen & Nguyen (2024). We construct equal-weighted quintile portfolios sorted on the moving targets score in a calendar-time framework, and we estimate Fama–MacBeth cross-sectional regressions with standard controls. This design allows us to test whether the LLM-based scores retain predictive power for subsequent stock returns, thereby providing a direct comparison to the original methodology.

### v3

@@ Abstract
Moving targets (cohen2024moving) – managers’ strategic shifting of key performance metrics when the original targets become difficult to achieve – have been shown to predict subsequent stock underperformance. However, our work reveals that the method employed in that study exhibits two key limitations that hinder the accuracy – noise in the extracted targets and loss of contextual information – both of which stem primarily from the use of a named entity recognition (NER). To address these two limitations, we propose an LLM-based target extraction method with a newly defined metric that better captures semantic context. This approach preserves semantic context beyond simple entity recognition and yields consistently higher predictive power than the original approach. Overall, our approach enhances the granularity and accuracy of financial text-based performance prediction.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have spurred efforts to leverage these tools for uncovering economically meaningful signals from unstructured financial texts such as earnings calls and corporate disclosures. In line with prior research, our work applies LLMs to enhance the target and metric extraction methodology proposed in cohen2024moving. Their study documents that managers strategically shift emphasis to new performance metrics when existing ones deteriorate – a “moving targets” behavior that predicts subsequent stock price declines.
@@ 2.1 Motivation
Prior work (cohen2024moving) has relied heavily on NER to extract performance targets from earnings call transcripts. However, we found that this approach suffers from limitations in both accuracy and explainability.
@@ 2.2 LLM-Based Target Extraction
Semantic Similarity. In the original formulation of the moving targets (cohen2024moving), the degree of movement is quantified as
@@
$MT_{c,t}=1-\frac{1}{n}\sum_{j=1}^{n}g(\tilde{S}_{j}).$
This semantic similarity–based metric extends the original discrete measure by preserving semantic equivalence even when targets are phrased differently. This metric quantifies the extent to which performance targets previously emphasized by a firm are dropped or semantically altered after four quarters, similar with cohen2024moving.
@@ 3.1 Data and Summary Statistics
We conduct our analysis using earnings call transcripts as the primary source of firm-level disclosures, following cohen2024moving. While we also experimented with 10-K and 10-Q MD&A sections, we found that earnings calls provided the most informative setting for detecting shifting in targets. Our study focuses on a restricted sample of S&P 100 firms due to computational cost. The sample covers between January 2010 and December 2024, producing 5,615 firm-quarter observations across 64 quarters.
@@ 3.2 Implementation Details
For validation, we adopt the empirical design of cohen2024moving. We construct equal-weighted quintile portfolios sorted on the moving targets score in a calendar-time framework, and we estimate Fama–MacBeth cross-sectional regressions with standard controls. This design allows us to test whether the LLM-based scores retain predictive power for subsequent stock returns, thereby providing a direct comparison to the original methodology.
@@ 4 Conclusion & Future Work
Figure 3: Representative example of targets that are extracted by NER-based method showing noise in the extracted targets, excerpted from Apple Inc. 2010 Q2 earnings call transcript. The NER-based method extracts generic terms such as year and units as targets illustrating the noisy nature of the extraction.
Figure 4: Representative example of targets that are extracted by NER-based method showing loss of contextual information, excerpted from NVIDIA Corporation 2025 Q2 earnings call transcript. The NER-based method extracts words like revenue as targets but drops surrounding qualifiers (e.g., Data Center revenue, Compute revenue), losing the specific type of revenue or operational driver. In addition, key product-related terms such as NVIDIA Hopper, and GPU computing are excluded from the target set even though the original approach is designed to include product entities, further underscoring the limitation of the NER-based extraction method.
Figure 5: Representative example of targets that are extracted by LLM-based method, excerpted from NVIDIA Corporation 2025 Q2 earnings call transcript. Here, we highlight both the phrase in the transcript where each target is detected and the corresponding extracted target itself. Unlike the NER-based approach, which often extracts overly generic terms such as year and units, the LLM-based method captures richer and more semantically precise business metrics (e.g., Quarterly Data Center revenue, Quarterly Compute revenue growth), reducing the noise in the extracted set.
Table 1: Statistics of the top 20 most frequent missing targets extracted by the NER-based method. The high frequency of generic phrases (e.g., the %, a % increase, the quarter) highlights the noise in the extracted targets, where context-free fragments dominate over concrete business metrics like specific revenue categories or operational drivers.
Rank Target Freq.
1 the % 4460
2 a % increase 3845
3 the range 3474
4 % growth 3456
5 q3 3006
6 q4 2853
7 the % range 2801
8 year 2716
9 the quarter 2499
10 share 2400
Rank Target Freq.
11 dollars 2268
12 a % 2146
13 earnings 2115
14 growth 2012
15 that % 1902
16 revenue 1882
17 a increase 1879
18 a range 1850
19 q1 1555
20 sales 1537
Table 2: Statistics of the top 20 most frequent missing targets extracted by the LLM-based method,. Unlike the NER-based approach—which often produces noisy, context-free fragments—our LLM-based extraction identifies substantive business metrics (e.g., market share, revenue growth, cash flow) and quantifies when these key targets are dropped or strategically shifted across consecutive year–quarters.
Rank Target Freq.
1 market share 706
2 share repurchases 589
3 revenue growth 582
4 pricing 562
5 cash flow 535
6 quarterly revenue growth 517
7 yearly guidance 484
8 free cash flow 480
9 dividend 458
10 quarterly earnings per share 455
Rank Target Freq.
11 quarterly share repurchases 441
12 quarterly revenue 436
13 profitability 435
14 dividends 416
15 gross margin 413
16 operating leverage 402
17 margins 402
18 revenue 388
19 share buybacks 377
20 margin expansion 368
Table 3: Calendar-time portfolio returns based on Moving Targets. We form equal-weighted quintile portfolios using the prior year’s distribution of Moving Targets across all stocks. Stocks enter portfolios in the month following each earnings conference call and are held until the next call. Panel A reports results using the original NER-based extraction method (cohen2024moving), while Panel B reports results using the LLM-based extraction method. We report Excess Returns (return minus risk-free rate), Fama–French 3-factor Alphas (market, size, and value), and 5-factor Alphas (market, size, value, momentum, and liquidity). $t$-statistics are in parentheses. , ∗∗, and ∗∗∗ indicate significance at the 10%, 5%, and 1% levels, respectively.
Panel A: NER-based extraction method
Q1 Q2 Q3 Q4 Q5 Q5–Q1
Excess Return 0.0156∗∗∗ 0.0131∗∗∗ 0.0111∗∗∗ 0.0132∗∗∗ 0.0135∗∗∗ -0.0031
(4.38) (3.85) (3.19) (3.99) (3.94) (-1.48)
3-Factor Alpha 0.0047∗∗∗ 0.0027∗∗ 0.0017 0.0040∗∗∗ 0.0039∗∗ -0.0018
(3.43) (2.12) (1.15) (2.88) (2.47) (-0.89)
5-Factor Alpha 0.0045∗∗∗ 0.0028∗∗ 0.0016 0.0040∗∗∗ 0.0041∗∗ -0.0014
(3.26) (2.19) (1.07) (2.79) (2.54) (-0.70)
Panel B: LLM-based extraction method
Q1 Q2 Q3 Q4 Q5 Q5–Q1
Excess Return 0.0151∗∗∗ 0.0135∗∗∗ 0.0123∗∗∗ 0.0128∗∗∗ 0.0122∗∗∗ -0.0039
(4.75) (3.94) (3.67) (3.50) (3.59) (-1.93)
3-Factor Alpha 0.0058∗∗∗ 0.0043∗∗∗ 0.0023∗∗ 0.0022 0.0021 -0.0047∗∗
(4.08) (3.06) (1.97) (1.44) (1.39) (-2.32)
5-Factor Alpha 0.0058∗∗∗ 0.0045∗∗∗ 0.0027∗∗ 0.0021 0.0017 -0.0050∗∗
(3.99) (3.10) (2.25) (1.36) (1.16) (-2.43)
Table 4: Fama–MacBeth cross-sectional regressions of individual firm-level stock returns on Moving Targets and standard return predictors. We compare results using the original NER-based extraction method (cohen2024moving) and the LLM-based extraction method. Log(Size) is the log of market value of equity, Log(BM) is the log of book value of equity over market value of equity, Ret(-1,0) is the previous month’s return, and Ret(-12,-1) is the cumulative return from month -12 to month -1. $t$-statistics are shown in parentheses. , ∗∗, and ∗∗∗ indicate significance at the 10%, 5%, and 1% levels, respectively.
(1) (2)
NER-based method LLM-based method
Ret
Moving Targets 0.0107 -0.0465
(1.10) (-1.70)
Log(Size) -0.0015 -0.0029
(-0.42) (-0.68)
Log(BM) -0.0016 -0.0006
(-0.36) (-0.17)
Ret(-1, 0) 0.0404 -0.0376
(0.98) (-0.71)
Ret(-12,-1) -0.0019 0.0029
(-0.22) (0.40)
Constant 0.0129 0.0828
(0.29) (1.43)
R-squared 0.3897 0.3884
N 5,615 5,615
Table 5: LLM prompt templates used in our method to extraction of targets.
Prompt <task> From the provided inputs, extract all performance metric and target labels and classify them into two separate lists according to the section in which they were mentioned: - presentation: items mentioned in the prepared presentation or opening remarks - analyst_qa: items mentioned during analyst question and answer sessions The term "performance metric and target labels" includes every explicitly stated financial result, non-financial result, and key business indicator discussed or mentioned. Both GAAP and non-GAAP labels if explicitly distinguished. ## Section classification - presentation: items mentioned in prepared remarks/opening statements before Q&A begins. Typical speakers: executives with "- Executives" in the name, before the first analyst question. - analyst_qa: items mentioned during the Q&A (analyst questions and executive answers after Q&A starts). Q&A typically starts after an Operator prompt like "[Operator Instructions]" or when an analyst first speaks (speaker label contains "- Analysts") and continues until the call ends. - Operator-only lines do not themselves contain items, but they can mark Q&A start. - Include only labels explicitly stated in content - Deduplicate near-duplicates to one normalized target per section </task> <inputs>earnings-call transcript as indexed JSON dialog</inputs> <output> - target: - Short noun phrase only - **Quarter results:** prefix with "Quarterly" and do NOT include any years or quarter numbers - **Fiscal-year results:** prefix with "Yearly" and do NOT include any years - **Guidance/targets/other indicators:** no period prefixes unless inherently part of the label; avoid dates, years, and quarter numbers - No numbers, units, currency symbols, or percent signs anywhere - Prefer consistent terminology across products/segments - index: integer index of the utterance containing the quote </output> <format> {{ "presentation": List[Dict], "analyst_qa": List[Dict] }} The dictionary must follow the following structure: {{ "target": str # Normalized short label for the metric/target according to the rules "index": int # Zero-based index of the utterance in the transcript where the quote appears }} </format>